SOME COMPLEXES OF METAL(II)CLORIDE WITH 3-C_{7}H_{9}N AND 4-C_{7}H_{9}N: PREPARATION, STRUC-TURAL CHARACTERISATION, MAGNETIC PROPERTIES, AND FT-IR AND ELECTRONIC SPECTRAL STUDIES

\underline{MUSTAFA KUMRU}, Fatih University, Faculty of Arts and Sciences, Department of Physics, 34900 Büyükcekmece, Istanbul, Turkey.